# A Heptalemma for Quantum Mechanics

### v1

@@ 4.6. Rejecting non-fragmentation
Fragmentalist views have been more readily adopted in the foundations of physics so far. Within the Copenhagen family of interpretations, Bohr’s complementarity pushes us in this direction. The quantum logic tradition (Wilce 2024) postulates that facts form an orthomodular lattice, but not a Boolean algebra. Bub and Pitowsky (2010) similarly point to non-Boolean event structures, emphasizing that while there is one objective world, it does not support a classical, global fact structure. Abramsky and Brandenburger (2011) provide a formalization of a fragmentalist reading of contextuality with sheaf theory.
@@ 5.1. Copenhagen
Strictly speaking, there is no unique Copenhagen interpretation, but only a family of overlapping views with a shared intellectual history going back to work done in Copenhagen in the 1920s (Faye 2024). When one speaks of the Copenhagen interpretation, one could be thinking of the views of any of a substantial list of physicists including Bohr, Heisenberg, Born, Pauli, Peierls, Landau, Rosenfeld, von Weizsäcker, and Peres. One could also have in mind a somewhat vague interpretational gloss that is broadly inspired by these and other authors’ ideas. This makes “the” Copenhagen interpretation harder to pin down in our heptalemma as well. Notwithstanding this diversity, it remains a coherent enough family to robustly reject Bell’s realism. Moreover, any set of ideas that is recognizably Copenhagen will certainly wish to preserve locality, measurement independence, and non-solipsism.
@@ 5.2. Everett
The Everett interpretation is popularly known as “the many-worlds interpretation” and was named “the relative state formulation” by its inventor. From the perspective of the heptalemma, however, it is neither a many-worlds interpretation nor a relational one. Specifically, under our definitions, the Everett interpretation upholds both one world and non-relationalism. The supposed “many worlds” of the Everett interpretation are just branches of the universal wavefunction evolution, and this straightforwardly fits into a single world, albeit one that is richer than conventionally assumed (namely, branching-tree-like). Chalmers (1996, p. 347) describes the Everett interpretation as a “one-big-world interpretation”: “There is only one world, but it has more in it than we might have thought” (1996, p. 347). Similarly, while speaking of a particular measurement outcome is only meaningful relative to the branch in which it occurs, the ontology of this view is that all possible outcomes actually do occur. Although we find ourselves in one branch, the reality is that all branches are equally real. All of these outcomes factually obtain, simpliciter, in a single coherent world. The label “multiverse” arguably describes this idea reasonably well, as long as it is understood that the multiverse in its entirety is a single world in the philosophical sense, not many.
@@
In the taxonomy of our heptalemma, we argue that the Everettians reject measurement realism. Particularly, they reject it in the second way discussed above. Measurement outcomes are facts for the Everettians; indeed, all possible measurement outcomes are. But this means that there is no single outcome for which one can have well-defined probabilities. Indeed, probabilities are not properly part of Everettianism. What is real in this view is the universal wavefunction and its unitary evolution. As in Quantum Darwinism, measurement is an add-on, which is – strictly speaking – not part of the physics except insofar as one should address how measurement can be understood to be part of the physical dynamics. Making contact with probabilities and with the Born rule is a research project for Everettians and has been approached from a variety of perspectives including decision theory (Deutsch 1999, Wallace 2012) and so-called self-locating probability (Saunders 1998, Sebens and Carroll 2018).
@@ 5.3. QBism
QBism starts out by taking a radically anti-realist stance towards the objects of quantum theory. At its most basic, it treats the quantum formalism as an aid to an agent’s decision-making. This sharply contrasts with interpretations that suppose that the objects of quantum mechanics have a directly physical or ontic character. Consistent with its early name “Quantum Bayesianism”, QBism asserts that quantum states and measurements do not represent anything that transcends the user’s unique situation, but that they capture the user’s capacity for action, physical embodiment, and any beliefs the user holds. On this picture, the objects of quantum mechanics are personal judgments that help the user to navigate the consequences of his or her actions (where a consequence is understood to be one’s personal experience of a measurement outcome). Moreover, QBists take themselves to interpret quantum theory normatively, namely as providing checks on the consistency of an agent’s beliefs, similar to checking whether a set of credences form a coherent set of probabilistic beliefs.
@@
We noted above how QBism responds to accusations of solipsism. Moreover, while it denies talk of hidden variables, QBism accepts measurement independence insofar as it affirms that the quantum-mechanical formalism can be used to help agents manage the consequences of their freely chosen actions. Locality, meanwhile, becomes almost tautological because, in QBism, the quantum formalism never describes physical influences between distant events; it is a tool for an agent to manage her own possible future experiences, and these always occur along her worldline, never at spacelike separations. [^11] In this context, the equivalent of non-locality would have to be detected within the agent’s mesh of beliefs in the form of a lack of independence of her probability assignments for the outcomes of her future actions on a distant system from her freely chosen actions on a nearby system.
@@
But what about the other three theses that jointly entail Bell’s realism? Does QBism affirm measurement realism as we have stated it in the heptalemma? While there is some subtlety, we think QBism is not only consistent with measurement realism, but should be read as robustly affirming it. There is no question that it regards measurement outcomes to be factual in the sense that outcomes are something that is the case. Outcomes are intrinsically personal, but they are not illusions, relational, or approximate notions. The facts in question may be first-personal, but they are facts no less. Indeed, QBism looks to measurement for a grounding of the notion of participatory realism (Fuchs 2017b). A recent paper further foregrounds QBism’s measurement realism, and situates it as pursuing the diametric opposite of most approaches to the quantum measurement problem, by deriving quantum dynamics from consistent reasoning about measurement (DeBrota et al. 2024). In other words, rather than somehow part of the dynamics, measurement is the most real thing to a QBist.
The subtlety comes from the second sense in which measurement realism can be rejected, which corresponds to the denial that there are well-defined probabilities for outcomes. QBism holds that there are well-defined probabilities, but it treats them, like measurement outcomes, as personal or subjective. Indeed, QBists view probabilities as degrees of belief for the consequences of one’s actions and even espouse a particularly radical form of subjective Bayesianism that identifies probabilities with betting dispositions. It is here that QBism might seem most anti-realist.
@@
If both of these are plausible readings, then, for now, there are two possible QBisms. The first of them, which we may label “Pluriverse QBism” [^13] It is ironic that the word “multiverse” is already associated with the Everett interpretation, as that interpretation manifestly affirms the one world thesis. In our treatment, if any interpretation deserved the term “many worlds”, it would be the version of QBism that rejects one world. On the other hand, “pluriverse” is uniquely appropriate for QBism due to its intellectual debts to William James., rejects one world, considering each agent to inhabit genuinely distinct subjective worlds that overlap in suitable ways. The second version of QBism, which may label “Fragmentalist QBism”, locates all observers in a single world, fundamentally fragmented along suitable notions of personalist fault lines. Pluriverse QBism might be easier to develop in some ways, as it can probably leverage more from existing philosophical literature. On the other hand, Fragmentalist QBism might be relatively more at home among the variety of fragmentalist constructions that have appeared in physics. Or, finally, it might turn out that one or the other QBism is ultimately incompatible with other convictions once considered closely enough.

### v2

@@ 4.6. Rejecting non-fragmentation
Fragmentalist views have been more readily adopted in the foundations of physics so far. Within the Copenhagen family of interpretations, Bohr’s complementarity pushes us in this direction. The quantum logic tradition (Wilce 2024) postulates that facts form an orthomodular lattice, but not a Boolean algebra. Bub and Pitowsky (2010) similarly point to non-Boolean event structures, emphasizing that while there is one objective world, it does not support a classical, global fact structure. With less emphasis on modifications to logic, the consistent histories approach (Griffiths 2024) proposes that the principal distinction of quantum mechanics lies in the availability of multiple mutually-incompatible modes of description, so-called frameworks, such that no single globally coherent totality of facts can obtain. Abramsky and Brandenburger (2011) provide a formalization of a fragmentalist reading of contextuality with sheaf theory.
@@ 5.1. Copenhagen
Strictly speaking, there is no unique Copenhagen interpretation, but only a family of overlapping views with a shared intellectual history going back to work done in Copenhagen in the 1920s (Faye 2024). When one speaks of the Copenhagen interpretation, one could be thinking of the views of any of a substantial list of physicists including Bohr, Heisenberg, Born, Pauli, Peierls, Landau, Rosenfeld, von Weizsäcker, and Peres. One could also have in mind a somewhat vague interpretational gloss that is broadly inspired by these and other authors’ ideas. This makes “the” Copenhagen interpretation harder to pin down in our heptalemma as well. Notwithstanding this diversity, it remains a coherent enough family to robustly reject Bell’s realism. Moreover, any set of ideas that is recognizably Copenhagen will certainly wish to preserve measurement independence and non-solipsism. Preservation of locality is almost universal as well, although there are some exceptions, notably Peres (1993).
@@ 5.2. Everett
The Everett interpretation is popularly known as “the many-worlds interpretation” and was named “the relative state formulation” by its inventor. From the perspective of the heptalemma, however, we suggest it should be viewed neither as a many-worlds interpretation nor as a relational one. Rather, it arguably upholds both one world and non-relationalism. The supposed “many worlds” of the Everett interpretation are just branches of the universal wavefunction evolution, and this straightforwardly fits into a single world, albeit one that is richer than conventionally assumed (namely, branching-tree-like). Chalmers (1996, p. 347) describes the Everett interpretation as a “one-big-world interpretation”: “There is only one world, but it has more in it than we might have thought” (1996, p. 347). Similarly, while speaking of a particular measurement outcome is only meaningful relative to the branch in which it occurs, the ontology of this view is that all possible outcomes actually do occur. Although we find ourselves in one branch, all branches are equally real. All of these outcomes factually obtain, simpliciter, in one coherent world. The label “multiverse” arguably describes this idea reasonably well, as long as it is understood that the multiverse in its entirety is a single world in the philosophical sense, not many.
@@
In the taxonomy of our heptalemma, we think the Everettians are best understood as rejecting measurement realism. Since all possible outcomes equally occur in the Everettian multiverse, albeit in different branches, there isn’t a fact of the matter as to which particular outcome actually results from a given measurement. At most, Everettians could say that all of those outcomes are facts. But then there is no single outcome for which one can have well-defined probabilities. Indeed, probabilities are not straightforwardly part of Everettianism. On this picture, Everettians reject measurement realism in the second way discussed above. What is real according to the Everett interpretation is the universal wavefunction and its unitary evolution. As in Quantum Darwinism, measurement is an add-on, which is – strictly speaking – not part of the physics except insofar as one should address how measurement can be understood to be part of the physical dynamics. Making contact with probabilities and with the Born rule is a research project for Everettians and has been approached from a variety of perspectives including decision theory (Deutsch 1999, Wallace 2012) and so-called self-locating probability (Saunders 1998, Sebens and Carroll 2018).
Could one also squeeze the Everett interpretation into different horns of our heptalemma, by understanding it in a slightly different way? One might say, for instance, that it gives up non-relationalism by making facts relative to branches. But the Everett interpretation is not normally understood as being committed to an ontology of dyadic facts in the way RQM is. Alternatively, one might say that the Everett interpretation postulates branch-specific facts, which are best viewed as indexical facts, akin to facts about where I am currently located or what time it is now. Assuming I find myself in a particular branch, it is an indexical fact that my measurement outcome is such and such. One might then say that each branch corresponds to a different total collection of indexical facts: a different “indexical world”. The Everett interpretation could thus be viewed as rejecting one world, after all, by postulating many indexical worlds. However, we suspect most Everettians would not want to “reify” those indexical worlds or view them as primitives of their physical ontology. Indeed, as Wallace (2012) has argued, the multiverse is an emergent phenomenon; it is not ontologically fundamental.
@@ 5.3. QBism
QBism starts out by taking a radically anti-objectivist stance towards quantum theory. At its most basic, it treats the quantum formalism as an aid to an agent’s decision-making. This sharply contrasts with interpretations that suppose that quantum mechanics seeks to describe the objective world. QBism asserts that quantum states and measurements do not represent anything that transcends the user’s unique situation, but that they capture the user’s capacity for action, physical embodiment, and any beliefs the user holds. On this picture, the objects of quantum mechanics are personal judgments that help the user to navigate the consequences of his or her actions (where a consequence is understood to be one’s personal experience of a measurement outcome). Moreover, QBists take themselves to interpret quantum theory normatively, namely as providing checks on the consistency of an agent’s beliefs, similar to checking whether a set of credences form a coherent set of probabilistic beliefs.
@@
We noted above how QBism responds to accusations of solipsism. Only a caricatured radical single-user version of QBism would be solipsistic. Moreover, while it denies talk of hidden variables, QBism accepts measurement independence insofar as it affirms that the quantum-mechanical formalism can be used to help agents manage the consequences of their freely chosen actions. Locality, meanwhile, becomes almost tautological because, in QBism, the quantum formalism never describes physical influences between distant events; it is a tool for an agent to manage her own possible future experiences, and these always occur along her worldline, never at spacelike separations. [^12] In this context, the equivalent of non-locality would have to be detected within the agent’s mesh of beliefs in the form of a lack of independence of her probability assignments for the outcomes of her future actions on a distant system from her freely chosen actions on a nearby system.
@@
But what about the other three theses that jointly entail Bell’s realism? Does QBism affirm measurement realism as we have stated it in the heptalemma? While there is some subtlety, we think QBism is not only consistent with measurement realism, but should be read as robustly affirming it. There is no doubt that it regards measurement outcomes as factual in the sense that outcomes are something that is the case. Outcomes are intrinsically personal, but they are not illusions, relational, or approximate notions. The facts in question may be first-personal, but they are facts no less. [^14] Importantly, they are of the usual monadic form: such and such is the case; they are not merely of the relational form: such and such is the case relative to such and such. Indeed, QBism looks to measurement for a grounding of the notion of participatory realism (Fuchs 2017b). A recent paper further foregrounds QBism’s measurement realism, and situates it as pursuing the diametric opposite of most approaches to the quantum measurement problem, by deriving quantum dynamics from consistent reasoning about measurement (DeBrota et al. 2024). In other words, rather than being somehow part of the dynamics, measurement is the most real thing according to QBism.
The subtlety comes from the second sense in which measurement realism can be rejected, which corresponds to the denial that there are well-defined probabilities for outcomes. QBism holds that there are well-defined probabilities, but it treats them, like measurement outcomes, as personal or subjective. QBists tend to view probabilities as degrees of belief for the consequences of one’s actions and even espouse a particularly radical form of subjective Bayesianism that identifies probabilities with betting dispositions. It is here that QBism might seem most anti-realist.
@@
If both of these are plausible readings, then, for now, there are two distinct versions of QBism, in addition to the implausible solipsistic version. The first version, which we may label “Pluriverse QBism” [^15] It is ironic that the word “multiverse” is already associated with the Everett interpretation, as that interpretation arguably affirms the one world thesis. In our treatment, if any interpretation deserved the term “many worlds”, it would be the version of QBism that rejects one world. On the other hand, “pluriverse” is uniquely appropriate for QBism due to its intellectual debts to William James., rejects one world and takes different agents to inhabit or experience genuinely distinct subjective worlds that overlap in suitable ways. The second version of QBism, which we may label “Fragmentalist QBism”, locates all observers in a single world, fundamentally fragmented along suitable notions of personalist fault lines.
There is a clear formal difference between these two versions of QBism. The formal machinery that would be needed to precisify each would look very different. The pluriverse version would postulate many subjective worlds and many subjective probability measures, where each of them individually behaves in the standard way, i.e., each subjective world is a complete and consistent collection of facts that hold at that world, and each subjective probability measure is a coherent assignment of probabilities to all elements of the relevant algebra of events. By contrast, the fragmentalist version would not postulate any coherent totalities at all. We would be dealing with a single world/reality that is an incoherent totality and of which only fragments are coherent. Moreover, we wouldn’t have a globally coherent probability assignment but only locally coherent subassignments. Pluriverse QBism might be easier to develop in some ways, as it can probably leverage more from existing philosophical literature. On the other hand, Fragmentalist QBism might be relatively more at home among the variety of fragmentalist constructions that have appeared in physics. Or, finally, it might turn out that one or the other version of QBism is ultimately incompatible with other convictions once considered closely enough.